OpenStudy (anonymous):

Find the difference quotient f(x+h)-f(x)/h , where h is not equal to zero for the function below: f(x)=5X^2 +3X Simplify the answer as much as possible.

OpenStudy (anonymous):

bunch of algebra ready?

OpenStudy (anonymous):

oh Yes. I have been trying to work it out but it is confusing.

OpenStudy (anonymous):

ok we have \(f(x)=5x^2+3x\) so the first thing we need to compute it \[f(x+h)=5(x+h)^2+3(x+h)\] by "compute" i really mean remove the parentheses

OpenStudy (anonymous):

make sure to square carefully, and don't forget to distribute the \(3\) \[f(x+h)=5(x+h)^2-3(x+h)\] \[f(x+h)=5(x+h)(x+h)+3(x+h)\] \[f(x+h)=5(x^2+2xh+h^2)+3x+3h\] and finally \[f(x+h)=5x^2+10xh+5h^2+3x+3x\] so far so good?

OpenStudy (anonymous):

typo on the last line, should have been \[f(x+h)=5x^2+10xh+5h^2+3x+3h\]

OpenStudy (anonymous):

you ok to this step? we are not done yet

OpenStudy (anonymous):

So far makes perfect sense

OpenStudy (anonymous):

from this we have to subtract \(f(x)\) again more algebra, don't forget to distribute the minus sign \[f(x+h)-f(x)=5x^2+10xh+5h^2+3x+3h-(5x^2+3x)\] \[f(x+h)-f(x)=f(x+h)=5x^2+10xh+5h^2+3x+3h-5x^2-3x\]

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

another damn typo \[f(x+h)-f(x)=5x^2+10xh+5h^2+3x+3h-5x^2-3x\]

OpenStudy (anonymous):

when you subtract, all terms without an \(h\) in them cancel (add up to zero) leaving only \[f(x+h)-f(x)=10xh+3h-5h^2\]

OpenStudy (anonymous):

finally you have to divide everything by \(h\) since there is an \(h\) in each term, divide them one by one making \[\frac{f(x+h)-f(x)}{h}=10x+3-5h\]

OpenStudy (anonymous):

That would be 10xh+3h-5h^2/h?

OpenStudy (anonymous):

yes, but when you divide by \(h\) divide each term since they all have an \(h\) in them, and you will end up with no denominator

OpenStudy (anonymous):

\[\frac{10xh+3h-5h^2}{h}=10x+3-5h\]

OpenStudy (anonymous):

ok I understand

OpenStudy (anonymous):

Thanks so much for your help you explained it very clearly.

OpenStudy (anonymous):

@satellite73 shouldn't \[5x^2+10xh+5h^2+3x+3h-5x^2-3x simplify \to 10xh+5h^2+3h\]?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes, did i not write that?

OpenStudy (anonymous):

oh damn, i had a minus sign there for no good reason you are right

OpenStudy (anonymous):

ok. Just making sure. So my final answer is \[10x+5h+3\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks again for your guidance.